OpenStudy (kweenlexis):

What value of x makes the equation true? –(9 + (–17)) = –9 + x

OpenStudy (kweenlexis):

is it -17???

OpenStudy (zpupster):

-(9-17) = -9 + x -9+17 =- 9+x add 9 both sides 17 = x run a check plug in for x -9+17 = -9+17

OpenStudy (kweenlexis):

so i was right?

OpenStudy (kweenlexis):

@GrammaticalMistake is 17 thr correct answer?

OpenStudy (kweenlexis):

the*

OpenStudy (anonymous):

Yep!

OpenStudy (kweenlexis):

thak you! i made a 100%

OpenStudy (anonymous):

Yay! :D

OpenStudy (kweenlexis):

lol xD